OpenStudy (anonymous):

Jennie purchased 5 of the 200 tickets at the raffle. There will be two drawings for a prize. The probability that Jennie will win both of the prizes is 1 out of ____. 1,592 1,600 1,990

ganeshie8 (ganeshie8):

in the first draw, she has a probability of 5 out of 200, or 1 out of 40 in the second draw, she has the same probability. so overall she has a probability of : \[\frac{1}{40}*\frac{1}{40}\] = \[\frac{1}{1600}\] that is 1 out of what ?

OpenStudy (anonymous):

1st time the prob is 5/200 ...second time it 4/199 multiply both of them....u get 1/ 1990

OpenStudy (anonymous):

Thanks

ganeshie8 (ganeshie8):

yep.. prophet is correct:) thanks @THE_PROPHET